OpenStudy (anonymous):

find the probability of drawing two hearts from a deck of 52 cards and obtaining heart cards both times, given that a heart is drawn the first time

OpenStudy (anonymous):

The question is not well defined.

OpenStudy (anonymous):

it is a conditional probabilty question

OpenStudy (anonymous):

\[\frac{12}{51}\]as there are 12 hearts left and 51 cards left assuming you do not replace the first heart

OpenStudy (anonymous):

oh, but the answer given in my book is 17/60

OpenStudy (anonymous):

then it is wrong

OpenStudy (anonymous):

you know the first card drawn was a heart, that is the "given" part so assuming it is not replaced, then the answer is \(\frac{12}{51}\) on the other hand, if it is replaced the answer is \(\frac{13}{52}=\frac{1}{4}\)

OpenStudy (anonymous):

That is assuming that the probability of choosing the heart is not 'loaded'.

OpenStudy (anonymous):

ok.. how about this one: At a certain university, 75% of the 1st-year engineering students passed math, 85% passed chemistry, and 90% passed math or chemistry. A 1st-year enginnering student is selected at random. Determine the estimated probability that the student passed math, given that the student passed chemistry